Calculate the volume of hydrogen produced at s.t.p when 25g of zinc is added to excess dilute hydrochloric acid at 31 degrees celsius and 778mmhg pressure.
1
Expert's answer
2019-08-15T03:55:01-0400

0.3846 mol                    X mol                      

      Zn + 2HCl = ZnCl2 + H2

   1 mol                         1 mol


n(Zn) = 25g / 65g/mol = 0.3846 mol.


n(H2) = X = 0.3846 mol.

V(H2) = n(H2) × Vm = 0.3846 mol × 22.4 L/mol = 8.6150 L ≈ 8.62 L.